# Computing the Elementary Symmetric Polynomials in Positive Characteristics

### v2

@@ Results.
While this explanation can provide some intuition behind the claim, the proof does not follow this argument. Instead, we analyze the coefficients of $f$ and construct a polynomial that “witnesses” the condition. Specifically, it is used to show that if a certain set of multilinear monomials have nonzero coefficients in $f$, then there must be another multilinear monomial that is nonzero. But this approach is perhaps more ideal, in a sense, to the one described above, as it is then obvious why this applies in the border setting. Recall that a property is called “closed,” meaning it being satisfied in the non-border setting implies that it is satisfied in the border setting, if there is a metapolynomial, meaning a polynomial whose variables are seen as the coefficients of an input polynomial, that evaluates to zero if and only if the input polynomial satisfies the property. It is thus often advantageous to explicitly give such a metapolynomial when showing that a property is closed.
@@ 2 Order-$2$ zero space of the elementary symmetric polynomials
In this section, we will focus on extending the results on the formula complexity of the elementary symmetric polynomials from [CKSV22], as in we will prove Theorem 1.1 and Lemma 1.7. Specifically, we will show that these results extend when we consider fields of positive characteristic, which we do by analyzing the order-$2$ zero set of the elementary symmetric polynomials, as this is the only part of the proof that uses the characteristic of the field. Upon careful inspection of [CKSV22], we can state the main result we rely on in the following lemma. For the sake of completeness, we prove this lemma in the appendix in Section B.
Lemma 2.1 ([CKSV22]).
@@
Then, in [CKSV22], the proof is completed through the following claim.
Claim 2.2 ([MZ17], Lemma 12 of [LMP19], Lemma 5.2 of [CKSV22]).
@@
In particular, we have that $V_{2}(e_{d}^{n})=\bigcup_{I\in{[n]\choose d-2}}V(x_{i}\mid i\in I)$.
One should notice that a formula complexity lower bound of $\Omega(d(n-d))$ on $e_{d}^{n}$ when $\operatorname{char}(\mathbb{F})=0$ immediately follows from this. We will spend this section studying this result over fields of positive characteristic.
@@
$e_{2}^{5}(\alpha,\alpha,\alpha,\alpha,\alpha)={5\choose 2}\alpha^{2}=10\alpha^{2},\ \frac{\partial e_{2}^{5}}{\partial x_{i}}(\alpha,\alpha,\alpha,\alpha,\alpha)=e_{1}^{4}(\alpha,\alpha,\alpha,\alpha)=4\alpha.$
@@
Before getting to the proof of this result, we believe that it is useful to provide some motivation for where the proof stems from. Specifically, we consider the case of $V_{2}(e_{2}^{n})$ for some $n\geq 2$. Observe that we can write each $\frac{\partial e_{2}^{n}}{\partial x_{i}}(a)=\sum_{j\neq i}a_{j}$. Then, for $i\neq j$, $\sum_{k\neq i}a_{k}-\sum_{k\neq j}a_{k}=a_{i}-a_{j}$. We thus conclude that, if $a\in V_{2}(e_{2}^{n})$, then $a_{1}=\dots=a_{n}$. As we repeat this strategy for increasing values of $d$, we notice that this strategy allows us to reduce to smaller degrees, where we observe that $a\in V_{2}(e_{d}^{n})$ if an only if the components of $a$ can be separated into at most d1d-1 groups, where the values in each group are the same. We will now formalize this approach.
@@
We will consider $S_{k}=\{(a_{1},\ldots,a_{n})\in\mathbb{F}^{n}\mid|\{a_{1},\ldots,a_{n}\}|\leq k\}$, which is the set of all points whose number of distinct coordinates is at most $k$. We will then show that $V_{2}(e_{d}^{n})\subseteq S_{d-1}$. From this, the claim immediately follows, as $\dim S_{k}\leq k$.
@@
For $d\geq 0$ and $m>d$, let $\alpha_{0},\ldots,\alpha_{d}\in\mathbb{F}$ be such that $\alpha_{d}=1$. Then, we have that $a\in V$ implies that $(a_{1},\ldots,a_{m})\in S_{d}$, for
$V=V\left(\sum_{i=0}^{d}\alpha_{i}e_{i}^{m-1}(x_{j}\mid j\in I)\ \middle|\ I\in{[m]\choose m-1}\right)$.
We can apply the above claim for d1d-1 and set $\alpha_{d-1}=1$ and $\alpha_{d-2}=\dots=\alpha_{0}=0$ to prove the lemma. We can now focus on proving the claim, which we will do by induction on $d$. Notice that the case of $d=0$ is obvious, as $\alpha_{0}=1$, so $V=\emptyset$. Now, consider the claim for $d$, using the inductive hypotheses for d1d-1. Let $a=(a_{1},\ldots,a_{n})\in V$. Observe that
@@
Hence, we have that either $a_{1}=a_{m}$ or $\sum_{i=1}^{d}\alpha_{i}e_{i-1}(a_{2},\ldots,a_{m-1})=0$. Observe that we can easily extend this argument to work for any $a_{k}$, for $k\in[m-1]$.
Now, let $I=\{k\in[m-1]\mid a_{k}=a_{m}\}$. Without loss of generality, assume that $m-1]\setminus I=[\ell]$. Observe that if $\ell<d$, then the claim trivially follows, so we assume that $\ell\geq d$. Consider some $k\in[\ell]$, and, using the fact that $a_{i}=a_{m}$ for each $i\in I$, observe that
$\begin{split}0&=\sum_{i=0}^{d-1}\alpha_{i+1}e_{i}(a_{j}\mid j\in[m-1]\setminus\{k\})=\sum_{i=0}^{d-1}\alpha_{i+1}\sum_{j=0}^{i}{|I|\choose i-j}a_{m}^{i-j}e_{j}(a_{p}\mid p\in[\ell]\setminus\{k\})\\ &=\sum_{j=0}^{d-1}\left(\sum_{i=j}^{d-1}\alpha_{i+1}{|I|\choose i-j}a_{m}^{i-j}\right)e_{j}(a_{p}\mid p\in[\ell]\setminus\{k\})=\sum_{j=0}^{d-1}\alpha_{j}^{\prime}e_{j}(x_{p}\mid p\in[\ell]\setminus\{k\}),\end{split}$
where $\alpha_{i}^{\prime}\in\mathbb{F}[x_{k+1},\ldots,x_{n},]$. Observe that $\alpha_{d-1}^{\prime}=\alpha_{d}\cdot{|I|\choose 0}=1$. We then can use the inductive hypothesis to conclude that $(a_{1},\ldots,a_{\ell})\in S_{d-1}$. Because $a_{\ell+1}=\dots=a_{m}$, we conclude that $a\in S_{d}$. ∎
Although the following information is sufficient to extend the theorem to fields of arbitrary characteristic, one may be interested in how tight this bound is. Specifically, suppose that we have fixed some field characteristic and some number of variables, then what is the relationship between the value of $d$ and the dimension of the order-$2$ zeros. We have already shown a simple example where the d1d-1 bound is tight, but we will now show that there are more such examples where it is tight (there are actually infinitely many of them).
@@
For $\operatorname{char}(\mathbb{F})=p\neq 0$ and $d\geq 1$, there is an $n\in\mathbb{Z}_{\geq 1}$ such that $\dim V_{2}(e_{d}^{n})\geq d-1$.
@@
Suppose that we have already picked some $n\geq d$. Consider arbitrary $\beta_{1},\ldots,\beta_{d-1}\in\mathbb{F}$. We will consider the point $\alpha=(\alpha_{1},\ldots,\alpha_{n})\in\mathbb{F}^{n}$ defined by
@@
$e_{d}^{n}(\alpha)=\sum_{i=2}^{\min(d,n-d+2)}{n-d+2\choose i}\left(\sum_{I\in{[d-2]\choose d-i}}\prod_{j\in I}\beta_{j}\right)\beta_{d-1}^{i}.$
@@
$\begin{split}\frac{\partial e_{d}^{n}}{\partial x_{k}}(\alpha)&=e_{d-1}^{n-1}(\alpha_{1},\ldots,\alpha_{k-1},\alpha_{k+1},\ldots,\alpha_{n})\\ &=\sum_{i=2}^{\min(d-1,n-d+2)}{n-d+2\choose i}\left(\sum_{I\in{[d-2]\setminus\{k\}\choose d-1-i}}\prod_{j\in I}\beta_{j}\right)\beta_{d-1}^{i}.\end{split}$
@@
$\begin{split}\frac{\partial e_{d}^{n}}{\partial x_{k}}(\alpha)&=e_{d-1}^{n-1}(\alpha_{1},\ldots,\alpha_{k-1},\alpha_{k+1}\ldots,\alpha_{n})\\ &=\sum_{i=1}^{\min(d-1,n-d+1)}{n-d+1\choose i}\left(\sum_{I\in{[d-2]\choose d-1-i}}\prod_{j\in I}\beta_{j}\right)\beta_{d-1}^{i}.\end{split}$
It is thus sufficient to pick an $n\geq d$ such that ${n-d+2\choose i}\equiv 0\mod p$ for every $i\in[2,\min(d,n-d+2)]$ and ${n-d+1\choose i}\equiv 0\mod p$ for every $i\in[1,\min(d-1,n-d+1)]$.
To do this, we fix $n$ so that $n-d+1\geq d-1$ and is a power of $p$. The proof will then be completed using Lucas’s Theorem.
Theorem 2.7 (Lucas’s Theorem, see [Luc78], [Mes14]).
@@
Observe that this shows that all of the above binomial coefficients are zero mod $p$, completing the proof. ∎
In the other direction, we would like to be able to adapt the proof from characteristic zero fields to find some conditions where this proof works. Unfortunately, the original proof from [MZ17], [LMP19], and [CKSV22] relies on the value nd+1n-d+1 being nonzero over the field, which is not necessarily preserved when using induction. Fortunately, if we combine this original proof with the result of Lemma 2.4, we can show that d2d-2 is sometimes tight.
@@
We begin by partitioning $V_{2}(e_{d}^{n})$ into sets based on the locations of where coordinates are zero. Specifically, given an index set $I\subseteq[n]$, we set
$\mathcal{S}_{I}=\{a\in\mathbb{F}^{n}\mid a_{i}\neq 0\Leftrightarrow i\in S,\ \forall i\in[n]\}.$
Observe that $\mathbb{F}^{n}=\bigcup_{I\subseteq[n]}\mathcal{S}_{I}$. We will then show that, for each $I\subseteq[n]$, we have that $\dim(\mathcal{S}_{I}\cap V_{2}(e_{d}^{n}))\leq d-2$ (where dimension is defined in terms of the Zariski topology), which will complete the proof.
Consider an arbitrary $I\subseteq[n]$. First, observe that if |I|<d1|I|<d-1, we have that $\dim\mathcal{S}_{I}\leq d-2$, so we get the desired result. Now, we can assume that $|I|\geq d-1$. Let $a\in\mathcal{S}_{I}\cap V_{2}(e_{d}^{n})$. Then, we can apply (6) and (7) to say

### v3

@@ 1.2 Proof techniques
While this explanation can provide some intuition behind the claim, the proof does not follow this argument. Instead, we analyze the coefficients of $f$ and construct a polynomial that “witnesses” the condition. Specifically, it is used to show that, if a certain set of multilinear monomials have nonzero coefficients in $f$, then there must be another multilinear monomial that is nonzero. This approach is preferable to the one described above, as it is then obvious why this applies in the border setting. Recall that a property is called “closed,” meaning it being satisfied in the non-border setting implies that it is satisfied in the border setting, if there is a metapolynomial, meaning a polynomial whose variables are seen as the coefficients of an input polynomial, that evaluates to zero if and only if the input polynomial satisfies the property. It is thus often advantageous to explicitly give such a metapolynomial when showing that a property is closed.
Note 1.8.
In the proofs of some of the theorems, we will sometimes require our field to be algebraically closed. Because the main results from this paper are lower bounds, these results also apply to fields that are not algebraically closed fields, as every field can be extended into one that is algebraically closed.
1.3 Prior Work
The earliest-known motivation for studying the complexity of the elementary symmetric polynomials came from boolean circuit complexity, where circuits compute boolean functions using a graph labeled with and-gates, or-gates, and not-gates. When considering boolean formulas of constant depth, it was shown in [13] that the majority function had super-polynomial lower bounds, where the majority function is given by $\bigvee_{I\in{[n]\choose n/2}}\bigwedge_{i\in I}x_{i}$. Clearly, $e_{n/2}^{n}$ looks like an algebraic analogue of the majority function, so it was believed to be a good candidate for a polynomial that would have constant-depth formulas with super-polynomial lower bounds. Unfortunately, due to a construction of Ben-Or (see Theorem 3.1 of [29]), $e_{d}^{n}$ was shown to have $\Sigma\Pi\Sigma$ formulas of size $O(n^{2})$ (but the problem of super-polynomial $\Sigma\Pi\Sigma$ lower bounds has since been solved, see [24]). In [28] and [29], it was shown that this upper bound is tight for $\Sigma\Pi\Sigma$ formulas of certain elementary symmetric polynomials over fields of characteristic zero. This was further extended in [6] to be tight over general algebraic formulas, also over fields of characteristic zero. In the more restrictive model of homogeneous multilinear formulas, where each node computes a homogeneous, multilinear polynomial, [18] found super-polynomial lower bounds for certain elementary symmetric polynomials (the result of which is independent of field characteristic).
In [29], the elementary symmetric polynomials were used to define an algebraic computational model, which was called the symmetric model. It was defined by taking a series of linear polynomials $L_{1},\ldots,L_{m}\in\mathbb{F}[x_{1},\ldots,x_{n}]$ and a degree $d\leq m$ and defining the computation by $e_{d}^{m}(L_{1},\ldots,L_{m})$. [29] then showed that the symmetric model is universal over fields of characteristic zero, meaning that it can compute any (homogeneous) polynomial. This was done using Fischer’s identity (see [10] and [29]), which tells us that we can write any degree $d$ homogeneous polynomial $f\in\mathbb{F}[x_{1},\ldots,x_{n}]$ as $f=L_{1}^{d}+\dots+L_{m}^{d}$, for some linear polynomials $L_{1},\ldots,L_{m}\in\mathbb{F}[x_{1},\ldots,x_{n}]$, and it was then shown that polynomials written in this form can be expressed in the symmetric model. We observe that this only works over fields of characteristic zero, as there are many polynomials in positive characteristic that cannot be written in this way, as, for example, $(x+y)^{p}=x^{p}+y^{p}$ when $\operatorname{char}(\mathbb{F})=p$.
Then, in [22], the results around the symmetric model were used to study the border “affine Chow rank” [^4] This name was selected based on [9] of a polynomial. We say that the affine Chow rank of a polynomial $f\in\mathbb{F}[x_{1},\ldots,x_{n}]$ is the smallest $k$ such that $f$ can be computed by a $\Sigma^{[k]}\Pi\Sigma$ circuit. We mention that there are polynomials with affine Chow rank that is $\Omega(n)$ over any field (see Lemma 3.2 of [7]). Recall that, given a complexity measure, we define its border complexity by extending the underlying field to $\mathbb{F}(\epsilon)$, for some new variable $\epsilon$, and, instead of requiring our model to compute $f(x)$, we need only to compute $\epsilon^{N}f+\epsilon^{N+1}F(x,\epsilon)$ for some $N\geq 0$ and $F\in\mathbb{F}[x,\epsilon]$ (where we may sometimes write that $\epsilon^{N}\cdot f+\epsilon^{N+1}\cdot F(x,\epsilon)\simeq f$). Through the results of the symmetric model, [22] showed that, over fields of characteristic zero, every homogeneous polynomial has border affine Chow rank of at most two.
@@ 2 Order-$2$ zero space of the elementary symmetric polynomials
In this section, we will focus on extending the results on the formula complexity of the elementary symmetric polynomials from [6]; we will prove Theorem 1.1 and Lemma 1.7. Specifically, we will show that these results extend when we consider fields of positive characteristic, which we do by analyzing the order-$2$ zero set of the elementary symmetric polynomials, as this is the only part of the proof that uses the characteristic of the field. Upon careful inspection of [6], we can state the main result we rely on in the following lemma. For the sake of completeness, we prove this lemma in the appendix in Section B.
Lemma 2.1 ([6]).
@@
Then, in [6], the proof is completed through the following claim.
Claim 2.2 ([26], Lemma 12 of [23], Lemma 5.2 of [6]).
@@
In particular, we have that $V_{2}(e_{d}^{n})=\bigcup_{I\in{[n]\choose d-2}}\{a\in\mathbb{F}^{n}\mid\forall\ i\in I,a_{i}=0\}$.
One should notice that a formula complexity lower bound of $\Omega(d(n-d))$ on $e_{d}^{n}$ when $\operatorname{char}(\mathbb{F})=0$ immediately follows from this. We will spend the rest of this section studying this result over fields of positive characteristic.
@@
$e_{2}^{5}(\alpha,\alpha,\alpha,\alpha,\alpha)={5\choose 2}\alpha^{2}=10\alpha^{2}=0,\ \frac{\partial e_{2}^{5}}{\partial x_{i}}(\alpha,\alpha,\alpha,\alpha,\alpha)=e_{1}^{4}(\alpha,\alpha,\alpha,\alpha)=4\alpha=0.$
@@
Before getting to the proof of this result, we believe that it is useful to provide some motivation for where the proof stems from. Specifically, we consider the case of $V_{2}(e_{2}^{n})$ for some $n\geq 2$. Observe that we can write each $\frac{\partial e_{2}^{n}}{\partial x_{i}}(a)=\sum_{j\neq i}a_{j}$. Then, for $i\neq j$, $\sum_{k\neq i}a_{k}-\sum_{k\neq j}a_{k}=a_{i}-a_{j}$. We thus conclude that, if $a\in V_{2}(e_{2}^{n})$, then $a_{1}=\dots=a_{n}$. As we repeat this strategy for increasing values of $d$, we observe that $a\in V_{2}(e_{d}^{n})$ if an only if the indices of $a$ can be partitioned into d1d-1 sets, where all indices in the same set correspond to the same value. We will now formalize this approach.
@@
We will show that all points in $V_{2}(e_{d}^{n})$ have at most d1d-1 distinct values in their coordinates. To define this rigorously, we first define vector spaces corresponding to each partition of $[n]$. Recall that, given a set $S$, we say that $\pi\subseteq 2^{[n]}$ is a partition of $S$ of size $k$ if $|\pi|=k$, $\emptyset\notin\pi$, $\bigcup_{U\in\pi}U=S$, and, for every $U,U^{\prime}\in\pi$ such that $U\neq U^{\prime}$, we have that $U\cap U^{\prime}=\emptyset$. We say that the set of all such $\pi$ is $\mathcal{P}_{k}(S)$. Given a partition $\pi\in\mathcal{P}_{k}([n])$, we define the vector space $P_{\pi}$ formed by $\pi$ to be the set of all points such that, for every pair of indices that are in the same set in the partition, they also always have the same value in the space. Rigorously, we say that
$P_{\pi}=\{a\in\mathbb{F}^{n}\mid\forall\ S\in\pi,\ \forall\ i,j\in S,\ a_{i}=a_{j}\}.$ (4)
Then, we can define the set of all points that can be split into $k$ groups with the same value by
$S_{k}=\bigcup_{\pi\in\mathcal{P}_{k}([n])}P_{\pi}.$
We will then show that $V_{2}(e_{d}^{n})\subseteq S_{d-1}$. From this, the claim immediately follows, as $\dim S_{k}=k$.
@@
For $d\geq 0$ and $m>d$, let $\alpha_{0},\ldots,\alpha_{d}\in\mathbb{F}$ be such that $\alpha_{d}=1$. Then, we have that
$\left\{a\in\mathbb{F}^{m}\ \middle|\ \forall\ i\in[m],\ \sum_{k=0}^{d}\alpha_{k}e_{k}^{m-1}(x_{j}\mid j\neq i)=0\right\}\subseteq S_{d}.$
We can apply the above claim for d1d-1 and set $\alpha_{d-1}=1$ and $\alpha_{d-2}=\dots=\alpha_{0}=0$ to prove the lemma. We will now focus on proving the claim, which we will do by induction on $d$. Notice that the case of $d=0$ is obvious, as $\alpha_{0}=1$, so the set is empty. Now, consider the claim for $d$, using the inductive hypotheses for d1d-1. Let $a=(a_{1},\ldots,a_{m})$ be such that, for every $i\in[m]$, $\sum_{k=0}^{d}\alpha_{k}e_{k}^{m-1}(a_{j}\mid j\neq i)=0$. Observe that
@@
Hence, we have that either $a_{1}=a_{m}$ or $\sum_{i=1}^{d}\alpha_{i}e_{i-1}^{m-2}(a_{2},\ldots,a_{m-1})=0$. Observe that we can easily extend this argument to any pair of indices.
Now, let $I=\{k\in[m-1]\mid a_{k}=a_{m}\}$. Without loss of generality, assume that $m-1]\setminus I=[\ell]$. Observe that if $\ell<d$, then the claim trivially follows, so we assume that $\ell\geq d$. Consider some $k\in[\ell]$, and, using the fact that $a_{i}=a_{m}$ for each $i\in I$ and $a_{k}\neq a_{m}$, observe that
$\begin{split}0&=\sum_{i=0}^{d-1}\alpha_{i+1}e_{i}^{m-2}(a_{j}\mid j\in[m-1]\setminus\{k\})=\sum_{i=0}^{d-1}\alpha_{i+1}\sum_{j=0}^{i}{|I|\choose i-j}a_{m}^{i-j}e_{j}^{\ell-1}(a_{p}\mid p\in[\ell]\setminus\{k\})\\ &=\sum_{j=0}^{d-1}\left(\sum_{i=j}^{d-1}\alpha_{i+1}{|I|\choose i-j}a_{m}^{i-j}\right)e_{j}^{\ell-1}(a_{p}\mid p\in[\ell]\setminus\{k\})=\sum_{j=0}^{d-1}\alpha_{j}^{\prime}e_{j}^{\ell-1}(x_{p}\mid p\in[\ell]\setminus\{k\}),\end{split}$
where $\alpha_{i}^{\prime}\in\mathbb{F}$. Observe that $\alpha_{d-1}^{\prime}=\alpha_{d}\cdot{|I|\choose 0}=1$. We then can use the inductive hypothesis to conclude that $(a_{1},\ldots,a_{\ell})\in S_{d-1}$. Because $a_{\ell+1}=\dots=a_{m}$, we conclude that $a\in S_{d}$. ∎
With the above proof, we have now extended the results of [6] to fields of positive characteristic. But, it is still independently interesting to get an understanding of how tight this bound is. We have shown a specific example where this lower bound is tight, but it would be nice to generalize this example for different number of variables and field characteristics. We do this in the following proposition.
@@
Let $\operatorname{char}(\mathbb{F})=p\neq 0$ and $n,d\geq 1$. If $n-d+1\equiv 0\mod p^{\lceil\log_{p}d\rceil}$ and $n\geq 2d-1$, then $\dim V_{2}(e_{d}^{n})\geq d-1$.
@@
Consider arbitrary $\beta_{1},\ldots,\beta_{d-1}\in\mathbb{F}$. We will consider the point $\alpha=(\alpha_{1},\ldots,\alpha_{n})\in\mathbb{F}^{n}$ defined by
@@
$e_{d}^{n}(\alpha)=\sum_{i=2}^{\min(d,n-d+2)}{n-d+2\choose i}e_{d-i}^{d-2}(\beta_{1},\ldots,\beta_{d-2})\beta_{d-1}^{i}.$
@@
$\begin{split}\frac{\partial e_{d}^{n}}{\partial x_{k}}(\alpha)&=e_{d-1}^{n-1}(\alpha_{1},\ldots,\alpha_{k-1},\alpha_{k+1},\ldots,\alpha_{n})\\ &=\sum_{i=2}^{\min(d-1,n-d+2)}{n-d+2\choose i}e_{d-1-i}^{d-3}(\beta_{1},\ldots,\beta_{k-1},\beta_{k+1},\ldots,\beta_{d-2})\beta_{d-1}^{i}.\end{split}$
@@
$\begin{split}\frac{\partial e_{d}^{n}}{\partial x_{k}}(\alpha)&=e_{d-1}^{n-1}(\alpha_{1},\ldots,\alpha_{k-1},\alpha_{k+1}\ldots,\alpha_{n})\\ &=\sum_{i=1}^{\min(d-1,n-d+1)}{n-d+1\choose i}e_{d-1-i}^{d-2}(\beta_{1},\ldots,\beta_{d-2})\beta_{d-1}^{i}.\end{split}$
It is thus sufficient to show that ${n-d+2\choose i}\equiv 0\mod p$ for every $i\in[2,\min(d,n-d+2)]$ and ${n-d+1\choose i}\equiv 0\mod p$ for every $i\in[1,\min(d-1,n-d+1)]$. We will do this using Lucas’s Theorem.
Theorem 2.7 (Lucas’s Theorem, see [25], [27]).
@@
We will now explain how to apply Lucas’s Theorem. Let $N\in\mathbb{N}$ and $a_{0},\ldots,a_{N}\in\mathbb{N}$ be such that $n-d+1=\sum_{k=0}^{N}a_{k}\cdot p^{k}$. Notice, because $n-d+1\equiv 0\mod p^{\lceil\log_{p}d\rceil}$, we have that $a_{k}=0$ for every $k\in[0,\lceil\log_{p}d\rceil]$. Now, consider some $i\in[2,d-1]$, and we will write it as its base $p$ expansion $i=\sum_{k=0}^{N}b_{k}\cdot p^{k}$. By applying Lucas’s Theorem, notice that
${n-d+2\choose i}=\left(\prod_{k=1}^{N}{a_{k}\choose b_{k}}\right)\cdot{1\choose b_{0}}=\left(\prod_{k=\lceil\log_{p}d\rceil+1}^{N}{a_{k}\choose b_{k}}\right)\left(\prod_{k=1}^{\lceil\log_{p}d\rceil}{a_{k}\choose b_{k}}\right)\cdot{1\choose b_{0}}$.
Because $i\in[2,d-1]$, either there is a $j\in[2,\lceil\log_{p}d\rceil]$ such that $b_{j}>0$ or $b_{0}>1$. Hence, ${n-d+2\choose i}=0$.
Then, for the other case, consider $i\in[1,d-1]$, and, similarly to previous, we write $i=\sum_{k=0}^{N}b_{k}\cdot p^{k}$. Then, by Lucas’s Theorem, we have that
${n-d+1\choose i}=\left(\prod_{k=0}^{N}{a_{k}\choose b_{k}}\right)=\left(\prod_{k=\lceil\log_{p}d\rceil+1}^{N}{a_{k}\choose b_{k}}\right)\left(\prod_{k=0}^{\lceil\log_{p}d\rceil}{0\choose b_{k}}\right)$.
Then, because $i<d$, we conclude that there is a $j\in[0,\lceil\log_{p}d\rceil]$ such that $b_{j}>0$, so ${n-d+1\choose i}=0$. ∎
With this in mind, we turn our attention to the following question: can we characterize when the space has dimension d1d-1 versus when it has dimension d2d-2. We will not be able to do this, but we will make some progress towards this. First, we would like to adapt the proof for characteristic zero fields to arbitrary fields (with some conditions). Unfortunately, the issue with the original proof is that it relies on nd+1n-d+1 not being zero, but it uses an inductive argument that may not preserve the difference between $n$ and $d$. But it turns out, if we combine this original proof with the result of Lemma 2.4, we can show that d2d-2 is sometimes tight.
@@
Consider an arbitrary vector space $V_{I}$, for some $I\subseteq[n]$, which is defined by points where all coordinates in $I$ are zero, i.e.,
$V_{I}=\{a\in\mathbb{F}^{n}\mid\forall\ i\in I,\ a_{i}=0\}.$
Without loss of generality, we will assume that I=[nk+1,n]I=[n-k+1,n]. Then, we will show that $V_{2}(e_{d}^{n})\cap V_{I}\subseteq V_{2}(f_{d})$, where $f_{d}=e_{d-1}^{n-k}(x_{1},\ldots,x_{n-k})\in\mathbb{F}[x_{1},\ldots,x_{n}]$. Then, by Lemma 2.4, we know that $\dim V_{2}(f_{d})=\dim V_{2}(e_{d-1}^{n-k})\leq d-2$. Thus, by taking the union over all $V_{I}$, we get the bound we are interested in.
Now, let $a\in V_{2}(e_{d}^{n})$ be such that $a_{1},\ldots,a_{n-k}\neq 0$ and $a_{n-k+1}=\dots=a_{n}=0$. We can apply (7) and (8) to say